OpenStudy (anonymous):

Find constants a and b so that (5, −4)is the solution of the system. ax + by = 6 bx + ay = −3. Find a and b?

OpenStudy (anonymous):

Substitute by the point in both equations & then solve simultaneously. 5a - 4b = 6 -------------> (1) 5b - 4a = -3------------> (2) Multiply (1) by 4 & (2) by 5 (1) = 20a - 16b = 24 (2) = -20a + 25b = -15 Add (1) & (2) after modification 9b = 9 , b = 1 Substitute in any equation to get a, a= 2

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Always welcome. :)